A grocer wants to mix peanuts worth $4.90 per pound with hazelnuts worth $7.90 per pounds to get 30 pounds of mixture that sells for $6.70 per pound. How many pounds of each should the grocer use?
Answer by Boreal(15235)

x=pounds of P
30-x=pounds of H
4.90x+7.90(30-x)=30(6.70)=201
4.9x+237-7.9x=201
-3x=-36
x=12 pounds of peanuts or $58.80
30-x=18 pounds of hazelnuts or $142.20
They add to $201.